# inkscapelatex=false

### v1

@@ 2.2. Discrete-time attention model
Let $w\in\mathbb{N}$ be a design parameter. The weight matrices at the $k$-th layer of the transformer, $k\in\mathbb{N}$, are denoted by $Q(k)\in\mathcal{M}_{w\times(n+1)}(\mathbb{R})$, $K(k)\in\mathcal{M}_{w\times(n+1)}(\mathbb{R})$ and $V(k)\in\mathcal{M}_{w\times(n+1)}(\mathbb{R})$, and are typically known as the Query, Key, and Value [^2] In the introduction we used $U$ to refer to the value matrix; this difference is resolved in this section. matrices, respectively. The input to the $k$-th layer is denoted by $\mathbf{x}=({\rm x}_{1},\ldots,{\rm x}_{\ell})\in\mathcal{M}_{(n+1)\times\ell}(\mathbb{R})$ and the output $\boldsymbol{z}\in\mathcal{M}_{w\times\ell}(\mathbb{R})$ of the self-attention mechanism is given by:
@@
where $\exp(\cdot)$ denotes the entry-wise exponential (i.e., $[\exp(R)]_{ij}=e^{R_{ij}}$), and $D(k)\in\mathcal{M}_{\ell\times\ell}(\mathbb{R})$ is defined as:
$D(k)_{ij}=\begin{cases}\left(\sqrt{n+1}\,\sum_{l=1}^{\ell}\exp({\rm x}_{l}(k)^{\top}\,K(k)^{\top}\,Q(k)\,{\rm x}_{i}(k))\right)^{-1},&i=j,\\ 0&i\neq j,\end{cases}\qquad 1\leq i,j\leq\ell.$
@@
Practical transformer applications often distribute the computations of the self-attention mechanism through several parallel heads, leading to what is commonly known as multi-headed self-attention. In this case, each layer of the transformer has $h\in\mathbb{N}$ heads. To make explicit the dependence on the head, we write (4) as:
$\boldsymbol{z}_{\eta}(k)=V_{\eta}(k)\,\mathbf{x}(k)\,D_{\eta}(k)\,\exp\left(\mathbf{x}(k)^{\top}\,K_{\eta}(k)^{\top}\,Q_{\eta}(k)\,\mathbf{x}(k)\right),\qquad 1\leq\eta\leq h.$
@@
The outputs from all attention heads are added after being multiplied by certain weight matrices $W_{\eta}\in\mathcal{M}_{(n+1)\times w}(\mathbb{R})$, $1\leq\eta\leq\ell$. Then, the resulting sum is added to the input of the layer $\mathbf{x}(k)$, using what is often called a skip connection. Lastly, a normalization function is applied to ensure that the output is bounded. In this work, we consider functions that normalize each token of the transformer separately, which is known as layer normalization and was first proposed in [28] as opposed to batch normalization, which consists of normalizing the distribution of the summed inputs. Hence, the normalization function is of the form:
$\mathbf{N}:\mathcal{M}_{(n+1)\times\ell}(\mathbb{R})\mapsto\mathcal{M}_{(n+1)\times\ell}(\mathbb{R}),\quad\mathbf{x}=({\rm x}_{1},\ldots,{\rm x}_{\ell})\mapsto\mathbf{N}(\mathbf{x})=(N({\rm x}_{1}),\ldots,N({\rm x}_{\ell})),$
for some $N:\mathbb{R}^{n+1}\to\mathbb{R}^{n+1}$. As mentioned before, our simplified model does not have a feedforward layer. Therefore, the output of the $k$-th layer is given by:
$\mathbf{x}(k+1)$ $=$ $\mathbf{N}\left(\mathbf{x}(k)+\sum_{\eta=1}^{h}W_{\eta}(k)\mathbf{z}_{\eta}(k)\right)$
$=$ $\mathbf{N}\left(\mathbf{x}(k)+\sum_{\eta=1}^{h}W_{\eta}(k)\,V_{\eta}(k)\,\mathbf{x}(k)\,D_{\eta}(k)\,\exp\left(\mathbf{x}(k)^{\top}\,K_{\eta}(k)^{\top}\,Q_{\eta}(k)\,\mathbf{x}(k)\right)\right).$
@@
Similarly to [15], in the following we consider the normalization function $N=\pi_{W}$ given in (2), which projects each token to the ellipsoid $\mathcal{E}_{W}^{n}$. In practice, this projection has been used explicitly in some models such as [29]. For clarity, we utilize the symbol $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})$ for the tokens evolving on the ellipsoid (after this explicit choice of normalization). The previous discrete-time dynamical system thus reads:
(5) $\boxed{\mathbf{y}(k+1)=\boldsymbol{\pi}_{W}\left(\mathbf{y}(k)+\sum_{\eta=1}^{h}W_{\eta}(k)\,V_{\eta}(k)\,\mathbf{y}(k)\,D_{\eta}(k)\,\exp\left(\mathbf{y}(k)^{\top}\,K_{\eta}(k)^{\top}\,Q_{\eta}(k)\,\mathbf{y}(k)\right)\right).}$
More explicitly, the discrete dynamics of the $i$-th token, $1\leq i\leq\ell$, is given by:
(6) ${\rm y}_{i}(k+1)=\pi_{W}\left({\rm y}_{i}(k)+\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}W_{\eta}(k)\,V_{\eta}(k)\,D_{\eta}(k)_{ii}\,\exp\left({\rm y}_{j}(k)^{\top}\,K_{\eta}(k)^{\top}\,Q_{\eta}(k)\,{\rm y}_{i}(k)\right){\rm y}_{j}(k)\right).$
@@ 2.3. Continuous-time attention model
Let $Y\in\mathfrak{X}((\mathcal{E}_{W}^{n})^{\ell})$ be a vector field and denote its flow by $Y^{\tau}:(\mathcal{E}_{W}^{n})^{\ell}\to(\mathcal{E}_{W}^{n})^{\ell}$. Given a map $g:(\mathcal{E}_{W}^{n})^{\ell}\times\mathbb{R}\to\mathbb{R}_{0}^{+}$ we use the notation $g({\rm y},\tau)=O_{\rm y}(\tau^{2})$ to denote the existence of a constant $T\in\mathbb{R}^{+}$ and of a function $\sigma:(\mathcal{E}_{W}^{n})^{\ell}\to\mathbb{R}_{0}^{+}$ so that for every $\tau\in[0,T]$ and for every ${\rm y}\in(\mathcal{E}_{W}^{n})^{\ell}$ we have $g({\rm y},\tau)\leq\sigma(x)\tau^{2}$. A map $\phi:(\mathcal{E}_{W}^{n})^{\ell}\times\mathbb{R}\to(\mathcal{E}_{W}^{n})^{\ell}$ is a first order approximation to the flow $Y^{\tau}$ if $\mathbf{d}(Y^{\tau}({\rm y}),\phi({\rm y},\tau))=O_{\rm y}(\tau^{2})$ where $\mathbf{d}$ denotes the distance on $(\mathcal{E}_{W}^{n})^{\ell}$ induced by the Euclidean distance on $(\mathbb{R}^{n+1}_{0})^{\ell}$.
Using the concepts introduced in the previous paragraph, our objective is to construct a vector field $Y$ so that the map defined by the right-hand side of (5) is a first order approximation of $Y^{\tau}$. To that effect we write $V_{\eta}$ as $V_{\eta}=\tau V_{\eta}^{\prime}$ and compute $Y$ as the best linear approximation in $\tau$ of (5). For simplicity, we work with (6), instead of (5), as rewrite it as:
@@
$f_{k}(\mathbf{y})=\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}W_{\eta}(k)\,V_{\eta}(k)^{\prime}\,D_{\eta}(k)_{ii}\,\exp\left({\rm y}_{j}^{\top}\,K_{\eta}(k)^{\top}\,Q_{\eta}(k)\,{\rm y}_{i}\right){\rm y}_{j},\quad\mathbf{y}\in(\mathcal{E}_{W}^{n})^{\ell}.$
For each $1\leq i\leq\ell$, the best linear approximation in $\tau$ is given by:
@@
$\dot{{\rm y}}_{i}=T_{{\rm y}_{i}}\pi_{W}\cdot f_{t}({\rm y}_{i})=T_{{\rm y}_{i}}\pi_{W}\cdot\left(\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}W_{\eta}(t)\,V^{\prime}_{\eta}(t)\,D_{\eta}(t)_{ii}\,\exp\left({\rm y}_{j}^{\top}\,K_{\eta}(t)^{\top}\,Q_{\eta}(t)\,{\rm y}_{i}\right){\rm y}_{j}\right),$
for each $1\leq i\leq\ell$, $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{W}^{n})^{\ell}$ and $t\geq 0$.
@@
$U_{\eta}(t)=W_{\eta}(t)\,V^{\prime}_{\eta}(t)\in\mathcal{M}_{(n+1)\times(n+1)}(\mathbb{R}),\quad P_{\eta}(t)=Q_{\eta}(t)^{\top}\,K_{\eta}(t)\in\mathcal{M}_{(n+1)\times(n+1)}(\mathbb{R}),$
for each $1\leq\eta\leq h$ and $t\geq 0$. We still refer to the matrix $U_{\eta}(t)$ as the value matrix since it plays a similar role. Similarly, we define the functions $\alpha_{ij}^{\eta},Z_{i}^{\eta}:\mathbb{R}_{0}^{+}\times(\mathcal{E}_{W}^{n})^{\ell}\to\mathbb{R}$ by:
$\alpha_{ij}^{\eta}(t,\mathbf{y})=\frac{1}{Z_{i}^{\eta}(t,\mathbf{y})}\,\exp({\rm y}_{i}^{\top}\,P_{\eta}(t)\,{\rm y}_{j}),\quad Z_{i}^{\eta}(t,\mathbf{y})=D_{\eta}(t)_{ii}^{-1}=\sqrt{n+1}\,\sum_{j=1}^{\ell}\exp({\rm y}_{i}^{\top}\,P_{\eta}(t)\,{\rm y}_{j}),$
respectively, for each $1\leq i,j\leq\ell$, $1\leq\eta\leq h$, $t\geq 0$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{W}^{n})^{\ell}$. The matrix having $\alpha_{ij}^{\eta}$ as its $i$th row and $j$th column entry is usually called the attention matrix of head $\eta$.
With the notation just introduced, the dynamical system that describes the evolution of a transformer with $h$ heads and $\ell$ tokens evolving on the ellipsoid $\mathcal{E}_{W}^{n}$ is given by:
(7) $\boxed{\dot{\rm y}_{i}=T_{{\rm y}_{i}}\pi_{W}\cdot\left(\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}\alpha_{ij}^{\eta}(t,\mathbf{y})\,U_{\eta}(t)\,{\rm y}_{j}\right)=\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}\alpha_{ij}^{\eta}(t,\mathbf{y})\left(U_{\eta}(t)\,{\rm y}_{j}-{\rm y}_{i}^{\top}\,W\,U_{\eta}(t)\,{\rm y}_{j}\,{\rm y}_{i}\right),}$
for each $1\leq i\leq\ell$, $t\geq 0$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{W}^{n})^{\ell}$. Let us denote by $Y$ the vector field on $(\mathcal{E}_{W}^{n})^{\ell}$ defined by (7). It is simple to check, by using Taylor’s theorem to expand the flow of $Y$ in powers of $\tau$, that (6) is a first order approximation to the flow of $Y$. A similar approach can be employed to derive a continuous-time model incorporating the effect of feedforward layers and a simple computation reveals that such model is a vector field of the form $Y+W$ where $W$ is a vector field describing a transformer with no attention layers. We view the results in this paper as a first step towards the analysis of the more complex model $Y+W$ that we leave to future work. The experimental results in Section 7 suggest the behavior of the more complex model is qualitatively the same as the behavior of the model (7) studied in this paper.
@@ 3. Transformers as gradient vector fields
It was noted in [13] that the transformer dynamics can be regarded as a gradient vector field under certain assumptions. For the benefit of the readers we formally prove such observation in the slightly more general setting where $P$ is not the identity matrix.
We consider the particular case of (7) with a single head, $h=1$, identity value matrix, $U_{1}(t)=\mathbb{I}_{n+1}$, and $P_{1}(t)=P$ time-independent, positive definite, and symmetric. In this case, we pick $W=P$, i.e.:
(8) $\dot{\rm y}_{i}=T_{{\rm y}_{i}}\pi_{P}\cdot\left(\sum_{j=1}^{\ell}\alpha_{ij}(\mathbf{y})\,{\rm y}_{j}\right)=\sum_{j=1}^{\ell}\alpha_{ij}(\mathbf{y})\left({\rm y}_{j}-{\rm y}_{i}^{\top}\,P\,{\rm y}_{j}\,{\rm y}_{i}\right),$
for each $1\leq i\leq\ell$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{P}^{n})^{\ell}$, where:
$\alpha_{ij}(\mathbf{y})=\frac{1}{Z_{i}(\mathbf{y})}\,\exp({\rm y}_{i}^{\top}\,P\,{\rm y}_{j}),\qquad Z_{i}(\mathbf{y})=\sqrt{n+1}\,\sum_{j=1}^{\ell}\exp({\rm y}_{i}^{\top}\,P\,{\rm y}_{j}).$

### v2

@@ 2.2. Discrete-time attention model
Let $w\in\mathbb{N}$ be a design parameter. The weight matrices at the $k$-th layer of the transformer, $k\in\mathbb{N}$, are denoted by $Q(k)\in\mathcal{M}_{w\times(n+1)}(\mathbb{R})$, $K(k)\in\mathcal{M}_{w\times(n+1)}(\mathbb{R})$ and $V(k)\in\mathcal{M}_{w\times(n+1)}(\mathbb{R})$, and are typically known as the Query, Key, and Value [^2] In the introduction we used $U$ to refer to the value matrix; this difference is resolved in this section. matrices, respectively. The input to the $k$-th layer is denoted by $\mathbf{x}\in\mathcal{M}_{(n+1)\times\ell}(\mathbb{R})$ and the output $\boldsymbol{z}\in\mathcal{M}_{w\times\ell}(\mathbb{R})$ of the self-attention mechanism is given by:
@@
where $\exp(\cdot)$ denotes the entry-wise exponential (i.e., $\exp(R)]_{ij}=e^{R_{ij}}$), and $D(k)\in\mathcal{M}_{\ell\times\ell}(\mathbb{R})$ is defined as:
$D(k)_{ij}=\left(\sqrt{n+1}\sum_{l=1}^{\ell}\exp({\rm x}_{l}(k)^{\top}K(k)^{\top}Q(k){\rm x}_{i}(k))\right)^{-1},$
if $i=j$, and $D(k)_{ij}=0$ otherwise.
@@
Practical transformer applications often distribute the computations of the self-attention mechanism through several parallel heads, leading to what is commonly known as multi-headed self-attention. To make explicit the dependence on the head, we write Eq.˜3 as:
$\boldsymbol{z}_{\eta}(k)=V_{\eta}(k)\mathbf{x}(k)D_{\eta}(k)\exp\left(\mathbf{x}(k)^{\top}K_{\eta}(k)^{\top}Q_{\eta}(k)\mathbf{x}(k)\right),$
for each $1\leq\eta\leq h$.
@@
The outputs from all attention heads are added after being multiplied by certain weight matrices $W_{\eta}\in\mathcal{M}_{(n+1)\times w}(\mathbb{R})$, $1\leq\eta\leq\ell$. Then, the resulting sum is added to the input of the layer $\mathbf{x}(k)$, using what is often called a skip connection. Lastly, a normalization function is applied to ensure that the output is bounded. In this work, we consider functions that normalize each token of the transformer separately, which is known as layer normalization and was first proposed in [28]. Hence, the normalization function $\mathbf{N}:\mathcal{M}_{(n+1)\times\ell}(\mathbb{R})\to\mathcal{M}_{(n+1)\times\ell}(\mathbb{R})$ is of the form:
$\mathbf{x}=({\rm x}_{1},\ldots,{\rm x}_{\ell})\mapsto\mathbf{N}(\mathbf{x})=(N({\rm x}_{1}),\ldots,N({\rm x}_{\ell})),$
@@
for some $N:\mathbb{R}^{n+1}\to\mathbb{R}^{n+1}$. Similarly to [25], in the following we consider the normalization function $N=\pi_{W}$ introduced in Section˜2.1, which projects each token to the ellipsoid $\mathcal{E}_{W}^{n}$. In practice, this projection has been explicitly used in some models such as [29]. For clarity, we utilize the symbol $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})$ for the tokens evolving on the ellipsoid (after this explicit choice of normalization). The resulting discrete-time dynamical system reads:
(4) ${\rm y}_{i}(k+1)=$
$\pi_{W}\left({\rm y}_{i}(k)+\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}W_{\eta}(k)V_{\eta}(k)D_{\eta}(k)_{ii}\exp\left({\rm y}_{j}(k)^{\top}K_{\eta}(k)^{\top}Q_{\eta}(k){\rm y}_{i}(k)\right){\rm y}_{j}(k)\right),$
where $1\leq i\leq\ell$ indexes each token.
Remark 2.3 (Standard layer normalization).
The standard layer normalization utilized in most transformers is given by:
$N({\rm x})=\frac{1}{\sigma({\rm x})}({\rm x}-\mu({\rm x})\mathbf{1})\star\gamma+\beta,$
for each ${\rm x}=(x^{1},\ldots,x^{n+1})\in\mathbb{R}^{n+1}$. In the previous expression, $\mathbf{1}$ denotes the vector $(1,\ldots,1)\in\mathbb{R}^{n+1}$, $\star$ denotes the element-wise product of vectors, and $\gamma,\beta\in\mathbb{R}^{n+1}$ are the learned scale and shift, respectively. Similarly, $\mu({\rm x})$ and $\sigma({\rm x})$ denote the mean and standard deviation of $\rm x$}, respectively. Under this normalization, we have:
$|N({\rm x})-\beta|^{2}$ $=\frac{1}{\sigma({\rm x})^{2}}\sum_{\mu=1}^{n+1}(x^{\mu}-\mu({\rm x}))^{2}(\gamma^{\mu})^{2}$
$=\frac{n+1}{\sum_{\mu=1}^{n+1}\left(x^{\mu}-\mu({\rm x})\right)^{2}}\sum_{\mu=1}^{n+1}(x^{\mu}-\mu({\rm x}))^{2}(\gamma^{\mu})^{2},$
where we used the notation $\gamma=(\gamma^{1},\ldots,\gamma^{n+1})$. It is clear that, if $\gamma=\gamma_{0}\mathbf{1}$ for some $\gamma_{0}\neq 0$, then the tokens lie on the $n$-sphere of center $\beta$ and radius $(n+1)\gamma_{0}^{2}$. Therefore, the results in this paper also apply to this type of normalization. We conjecture that, for arbitrary scale parameters, the tokens will also lie on a certain hypersurface of $\mathbb{R}^{n+1}$ diffeomorphic to the $n$-sphere, but a more careful analysis has to be carried out to understand the geometry of such hypersurface.
@@ 2.3. Continuous-time attention model
In this section we introduce additional notation that is only used to derive the continuous time model. Readers not interested in the model’s derivation can skip the next two paragraphs and start reading the paragraph commencing with “To simplify notation”.
Let $Y\in\mathfrak{X}((\mathcal{E}_{W}^{n})^{\ell})$ be a vector field and denote its flow by $Y^{\tau}:(\mathcal{E}_{W}^{n})^{\ell}\to(\mathcal{E}_{W}^{n})^{\ell}$. Note that $Y$ is complete, i.e., its flow is defined for each $\tau\in\mathbb{R}$, since $(\mathcal{E}_{W}^{n})^{\ell}$ is compact. Given a map $g:(\mathcal{E}_{W}^{n})^{\ell}\times\mathbb{R}\to\mathbb{R}_{0}^{+}$, we use the notation $g(\mathbf{y},\tau)=O_{\mathbf{y}}(\tau^{2})$ to denote the existence of a constant $T\in\mathbb{R}^{+}$ and a function $\sigma:(\mathcal{E}_{W}^{n})^{\ell}\to\mathbb{R}_{0}^{+}$ such that, for each $\tau\in[0,T]$ and $\mathbf{y}\in(\mathcal{E}_{W}^{n})^{\ell}$, we have $g(\mathbf{y},\tau)\leq\sigma(\mathbf{y})\tau^{2}$. A map $\phi:(\mathcal{E}_{W}^{n})^{\ell}\times\mathbb{R}\to(\mathcal{E}_{W}^{n})^{\ell}$ is a first order approximation to the flow $Y^{\tau}$ if $\mathbf{d}(Y^{\tau}(\mathbf{y}),\phi(\mathbf{y},\tau))=O_{\mathbf{y}}(\tau^{2})$ where $\mathbf{d}$ denotes the distance on $(\mathcal{E}_{W}^{n})^{\ell}$ induced by the Euclidean distance on $(\mathbb{R}^{n+1}_{0})^{\ell}$.
Using the concepts introduced in the previous paragraph, our objective is to construct a vector field $Y$ such that the map defined by the right-hand side of Eq.˜4 is the best first order approximation of $Y^{\tau}$. To that end, we write $V_{\eta}(k)$ as $V_{\eta}(k)=\tau V_{\eta}^{\prime}(k)$ for each $1\leq\eta\leq h$, with $0<\tau\ll 1$ being a small parameter. Hence, Eq.˜4 may be rewritten as:
@@
$f_{k}(\mathbf{y})=$ $\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}W_{\eta}(k)V_{\eta}(k)^{\prime}D_{\eta}(k)_{ii}\exp\left({\rm y}_{j}^{\top}K_{\eta}(k)^{\top}Q_{\eta}(k){\rm y}_{i}\right){\rm y}_{j},$
for each $\mathbf{y}\in(\mathcal{E}_{W}^{n})^{\ell}$. For each $1\leq i\leq\ell$, the best linear approximation in $\tau$ is given by:
@@
(5) $\dot{{\rm y}}_{i}=T_{{\rm y}_{i}}\pi_{W}\cdot\left(\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}W_{\eta}(t)V^{\prime}_{\eta}(t)D_{\eta}(t)_{ii}\exp\left({\rm y}_{j}^{\top}K_{\eta}(t)^{\top}Q_{\eta}(t){\rm y}_{i}\right){\rm y}_{j}\right),$
with $1\leq i\leq\ell$, $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{W}^{n})^{\ell}$, and $t\in\mathbb{R}_{0}^{+}$, as the differential equation whose solution provides the best first order approximation of Eq.˜4.
@@
$U_{\eta}(t)=W_{\eta}(t)V_{\eta}^{\prime}(t)\in\mathcal{M}_{(n+1)\times(n+1)}(\mathbb{R}),$
$P_{\eta}(t)=Q_{\eta}(t)^{\top}K_{\eta}(t)\in\mathcal{M}_{(n+1)\times(n+1)}(\mathbb{R}),$
for each $1\leq\eta\leq h$ and $t\in\mathbb{R}_{0}^{+}$. We still refer to the matrix $U_{\eta}(t)$ as the value matrix since it plays a similar role. Similarly, we define the following functions:
$\alpha_{ij}^{\eta}:\mathbb{R}_{0}^{+}\times(\mathcal{E}_{W}^{n})^{\ell}\to\mathbb{R},\quad$ $\alpha_{ij}^{\eta}(t,\mathbf{y})=\frac{1}{Z_{i}^{\eta}(t,\mathbf{y})}\exp({\rm y}_{i}^{\top}P_{\eta}(t){\rm y}_{j}),$
$Z_{i}^{\eta}:\mathbb{R}_{0}^{+}\times(\mathcal{E}_{W}^{n})^{\ell}\to\mathbb{R},\quad$ $Z_{i}^{\eta}(t,\mathbf{y})=D_{\eta}(t)_{ii}^{-1}=\sqrt{n+1}\sum_{j=1}^{\ell}\exp({\rm y}_{i}^{\top}P_{\eta}(t){\rm y}_{j}),$
respectively, for each $1\leq i,j\leq\ell$, $1\leq\eta\leq h$, $t\in\mathbb{R}_{0}^{+}$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{W}^{n})^{\ell}$. The matrix having $\alpha_{ij}^{\eta}$ as its $i$-th row and $j$-th column entry is usually called the attention matrix of head $\eta$.
With the notation just introduced, the dynamical system that describes the evolution of $\ell$ tokens on the ellipsoid $\mathcal{E}_{W}^{n}$ according to a transformer with $h$ heads is given by:
(6) $\boxed{\dot{\rm y}_{i}=T_{{\rm y}_{i}}\pi_{W}\cdot\left(\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}\alpha_{ij}^{\eta}(t,\mathbf{y})U_{\eta}(t){\rm y}_{j}\right),}$
for each $1\leq i\leq\ell$, $t\in\mathbb{R}_{0}^{+}$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{W}^{n})^{\ell}$.
@@ 3. Transformers as gradient vector fields
It was noted in [30] that the transformer dynamics can be regarded as a gradient vector field under certain assumptions. For the benefit of the readers we formally prove such observation in the slightly more general setting where $P$ is not the identity matrix. Throughout this section, we consider the particular case of Eq.˜6 described by the following assumptions.
Hypothesis 3.1.
There is only one head, $h=1$, $W=P$ and we have:
  1. (1)
    $U_{1}(t)=\mathbb{I}_{n+1}$, and
  2. (2)
    $P_{1}(t)=P$ is time-independent, positive definite, and symmetric.